VIBRATIONAL SPECTROSCOPY AND CHEMICAL DYNAMICS IN REACTANT COMPLEXES OF HY-DROXYL RADICALS WITH HYDROGEN AND METHANE^{a}

\underline{MARSHA I. LESTER}, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323.

The bimolecular reactions of hydroxyl radicals with hydrogen and methane play a fundamental role in the combustion of hydrocarbon fuels. This laboratory is examining the OH + H_{2} and OH + CH_{4} hydrogen abstraction reactions from a new perspective by trapping the reactants within a shallow attractive well in the entrance channel to reaction. The resultant H_{2}-OH and CH_{4}-OH reactant complexes are then vibrationally activated through infrared or stimulated Raman excitation. A variety of double resonance schemes have been implemented to record the spectra, lifetime, and product state distributions of the vibrationally activated complexes. Vibrational spectroscopy provides an effective means to examine the structural parameters of these complexes in their ground electronic state. Intermolecular bending excitation is found to control the relative orientation of the reactants within the complex, producing some that resemble the transition state structure. The OH, H_{2}, or CH_{4} vibrational excitation also supplies sufficient energy to surmount the activation barrier and thereby may induce chemical reaction. Alternatively, the vibrational excitation can cause the weak intermolecular bond to break through vibrational predissociation. Time- and frequency-domain measurements reveal how long the initial excitation stays localized in the OH, H_{2}, or CH_{4} mode and how excess energy is partitioned among the products. The experimental results will also be compared with theoretical calculations of the observables to yield new insights on the potential energy surfaces for the OH + H_{2} and OH + CH_{4} systems.

aThe Office of Basic Energy Sciences of the Department of Energy has supported this research. The Chemistry Division of the National Science Foundation has also provided partial equipment support.